A random sample of 13 customers in a Kentucky fast food restaurant revealed an average bill of OMR 18.75 per person. The population standard deviation is OMR5.4. Estimate 98% confidence
interval for the mean bill of all customers.


A random sample of n customers in a Kentucky fast food restaurant revealed an average bill of OMR & per person. The popula
Answer #1

n = 13

T = 18.75

\sigma = 5.4

\alpha = 0.02

From Table, critical values of Z = \pm 2.326

Confidence Interval:

\bar{x}\pm \frac{Z\times \sigma }{\sqrt{n}}=18.75\pm \frac{2.326\times 5.4}{\sqrt{13}}=18.75\pm 3.484=(15.266,22.234)

So,

Answer is:

(15.266, 22.234)